OpenStudy (anonymous):

Solve using the quadratic formula. x² + x = 12

OpenStudy (saifoo.khan):

a = 1 b = 2 c = -12 Use these values and insert those in quadratic formula.

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + x = 12\] \[\Large x^2 + x - 12 = 0\] Now use the quadratic formula to solve for x \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(1)\pm\sqrt{(1)^2-4(1)(-12)}}{2(1)}\] \[\Large x = \frac{-1\pm\sqrt{1-(-48)}}{2}\] \[\Large x = \frac{-1\pm\sqrt{49}}{2}\] \[\Large x = \frac{-1+\sqrt{49}}{2} \ \text{or} \ x = \frac{-1-\sqrt{49}}{2}\] \[\Large x = \frac{-1+7}{2} \ \text{or} \ x = \frac{-1-7}{2}\] \[\Large x = \frac{6}{2} \ \text{or} \ x = \frac{-8}{2}\] \[\Large x = 3 \ \text{or} \ x = -4\] So the solutions are \[\Large x = 3 \ \text{or} \ x = -4\]

OpenStudy (anonymous):

@jim_thompson5910 best answer I've seen so far ! well done !

OpenStudy (anonymous):